OpenStudy (anonymous):

How do I find the domain and range in this quadratic: y=x^2-2x+4?

OpenStudy (anonymous):

For polynomials, the domain is always going to be \([-\infty, \infty]\)

OpenStudy (anonymous):

For quadratics, to find the range you find the vertex and figure out if it opens up or down.

OpenStudy (anonymous):

So how do I write the domain exactly ?

OpenStudy (anonymous):

\[ \begin{split} y &=x^2-2x+4 \\ &= (x-1)^2+3 \\ y-3 &= (x-1)^2 \end{split} \]This means the vertex is at \((1, 3)\)

OpenStudy (anonymous):

To write the domain you'd just write \((-\infty , \infty ) \)

OpenStudy (anonymous):

Or you could say "all real numbers"

OpenStudy (anonymous):

Oh How do you know it's all real nimbers?

OpenStudy (anonymous):

For these types of problems, we generally ignore complex numbers. The reason we know it is all real numbers is because we never divide by 0 or take the square root of a negative number.

OpenStudy (anonymous):

Okay, Thanks (:

OpenStudy (anonymous):

Is it always all real numbers for the domain ?

OpenStudy (anonymous):

For polynomials, yes. For a function like \(f(x)=1/x\), no.

OpenStudy (anonymous):

Would the range be y is -1 ?

OpenStudy (anonymous):

Or greater than or equal to or less than or equal to ?

OpenStudy (anonymous):

It's opening up, so it can't go any lower than the vertex.

OpenStudy (anonymous):

where did the \(-1\) come from?

OpenStudy (anonymous):

On one of my older papers they switched the signs.

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{a}\) in your example it is \[-\frac{-2}{2\times 1}=1\] the second coordinate is what you get when you replace \(x\) by \(1\), in this case you get \(y=1^2-2\times 1+4=3\)

OpenStudy (anonymous):

But what confusing me is I don't know how to write it. The range.

OpenStudy (anonymous):

since \(3\) is the lowest the curve can be, it is \([3,\infty)\)